OpenStudy (anonymous):

hello, can someone help me with this, please: Find all solutions of sin x + cos x - sin x cos x = sin^2x + cos^2x in the interval [0; 2]

OpenStudy (anonymous):

do you have some idea where to start?

OpenStudy (anonymous):

yep, simplifying, but kinda stuck there

OpenStudy (anonymous):

well, on that right side sin^2x+cos^2x equals 1 right

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

but i dont think that's the right way tho

OpenStudy (anonymous):

do you have any idea?

OpenStudy (anonymous):

you know i am trying to do some calculus work myself, and i am also trying to think of something for your problem, give me 5 min

jimthompson5910 (jim_thompson5910):

sin x + cos x - sin x cos x = sin^2x + cos^2x sin x + cos x - sin x cos x = 1 sin x + cos x - sin x cos x - 1 = 0 sin x - sin x cos x + cos x- 1 = 0 (sin x - sin x cos x) + (cos x- 1) = 0 -sin x(-1 + cos x) + 1(cos x- 1) = 0 -sin x(cos x -1) + 1(cos x- 1) = 0 (-sin x + 1)(cos x- 1) = 0 -(sin x - 1)(cos x- 1) = 0 (sin x - 1)(cos x- 1) = 0 From here, solve sin x - 1 = 0 and cos x - 1 = 0

OpenStudy (anonymous):

thank you so much